# DiffImp: Efficient Diffusion Model for Probabilistic Time Series Imputation with Bidirectional Mamba Backbone

### v1

@@ 4.6 Ablation Studies
Bidirectional Yes Yes 5.46e-4$\pm$1.6e-5 1.17e-2$\pm$7.4e-5 1.21$\pm$7.5e-32.33e-2$\pm$3.1e-4
Forward Yes Yes 7.19e-4$\pm$2.0e-5 1.26e-2$\pm$2.1e-4 1.29$\pm$2.2e-22.67e-2$\pm$2.9e-4
Forward Yes No 7.48e-4$\pm$9.5e-5 1.23e-2$\pm$3.5e-4 1.23$\pm$3.5e-22.71e-2$\pm$1.5e-3
Backward Yes Yes 7.24e-4$\pm$7.3e-5 1.30e-2$\pm$4.2e-4 1.30$\pm$4.2e-22.69e-2$\pm$1.3e-3
Backward Yes No 8.39e-4$\pm$6.1e-5 1.46e-2$\pm$3.8e-4 1.46$\pm$3.8e-22.89e-2$\pm$1.0e-3
Bidirectional Yes No 8.85e-4$\pm$2.8e-5 1.40e-2$\pm$3.1e-4 1.44$\pm$3.4e-22.97e-2$\pm$4.8e-4
Bidirectional No Yes 9.66e-4$\pm$9.5e-5 1.53e-2$\pm$3.3e-4 1.57$\pm$3.5e-23.09e-2$\pm$1.3e-3
5 Conclusion and Future Work
In this paper, we propose DiffImp, a time series imputation model based on DDPM and Mamba backbone, which incorporates bidirectional information flow, temporal attention and inter-variable dependencies. DiffImp enables efficient time series modeling with linear complexity. Experimental results demonstrate that DiffImp achieves superior performance across multiple datasets, various missing patterns, and different missing ratios.
For future work, one possible direction is to further reduce the time complexity of the sampling process while already lowering the complexity of time series modeling, in order to enhance the model’s inference efficiency. Another possible direction is to extend the application of diffusion models by applying DiffImp to other time series downstream tasks and time series representation learning tasks.
6 Reproducibility
To ensure reproducibility and facilitate experimentation, datasets and code are available at: https://anonymous.4open.science/r/DiffImp-843F.

### v2

@@ 4.2.3. Experimental Results and Analysis
[…]
As stated in Sec.3, our method employs a modern state space model, Mamba (Dao and Gu, 2024) as the backbone. Similar to other methods, we also adopt special designs to capture inter- and intra- channel dependencies. The results in Table.3 and 4 show that our method outperforms those with transformer and S4-based core architectures. This demonstrates: 1) the effectiveness of Mamba structure as the backbone for diffusion models and 2) the validity of the designs we adopted in addressing the problem of time series imputation.
4.3. Visualization Results
Figure 4. Visualization results of probabilistic time series imputation results on MuJoCo, AQI and Physionet dataset. The orange solid line represents the mean of the imputation results, while the dark green represents the 50% confidence interval, and the light green solid line represents the 95% confidence interval. In Figure a, the blue solid line represents the target of imputation (i.e., the ground truth). In Figures b and c, the blue dots represent the ground truth of the missing points, and the red crosses represent the observed values.
Fig.4 shows the visualization results on the MuJoCo dataset, AQI dataset and Physionet dataset. We can see that almost all ground truth values for the points to be imputed fall within the 95% confidence interval, and most of the ground truth values are within the 50% confidence interval, which demonstrates the effectiveness of our method.
4.4. Ablation Studies
Effectiveness of Proposed Modules. To validate the effectiveness of the proposed module, we conduct ablation experiments on the following aspects: 1) the bidirectional modeling, 2) the temporal attention mechanism, 3) the inter-channel multivariate dependencies. We also replace the CMB block with channel attention module implemented using (Hu et al., 2018) to validate the effectiveness of CMB block. All experiments are conducted on the MuJoCo dataset with the missing ratio 90%.
The results are shown in Table.5. It can be observed that the module equipped with BAM and CMB block performs the best, significantly outperforming the results of removing any one of these components across all four metrics. The temporal attention module has the largest impact on the model, and its removal leads to a significant performance drop. Similarly, removing the CMB module also results in a notable degradation in performance, meanwhile, modifying the CMB block to Channel Attention in (Hu et al., 2018) also leads to a performance drop, which proves that state space models are also effective for modeling inter-channel dependencies. On the other hand, adjusting the BAM module to its unidirectional form also causes some degree of performance decrease, which means the bidirectional modeling manner is capable of modeling more intra-channel dependency. This fully demonstrates the effectiveness of all our proposed blocks and the capability of linear state space models as backbones for time series diffusion models.
Time Efficiency of Mamba Backbones. We compare the inference time of our model with a variant using the transformer backbone. The results is shown in Table.6. We can see that our model achieves faster sampling, less memory cost comparing with the model with transformer backbones, which demonstrates the efficiency of our model and our proposed backbone.
Table 6. Inference time comparison with transformer backbones
Model Transformer-backbone SSD-TS(128)
#Parameters(M) 9.208 87.57
Inference time/sample(s) 1.5126 0.93
GPU Memory(MB) 11250 4536
4.5. Parameter Sensitivity Analysis
We conduct parameter sensitivity experiment and the results are shown in Table.7. There are three hyperparameters with different dimensions: sequence, residual connection, and input projection dimension, which are set to be equal in our experiments. We test different results for $C=32,64,128$. The experimental results show that as $C$ increases, all metrics significantly decrease. Additionally, since $C=256$ exceeds one single GPU memory capacity, we choose $C=128$ in our experiments to balance between metrics and computational cost.
Table 7. Parameter Sensitivity Results
#Channel MAE MSE RMSE
32 0.0482 0.0066 0.0809
64 0.0147 0.00075 0.0273
128 0.0135 0.00065 0.0254
5. Related Work
5.1. Time Series Imputation
The objective of time series imputation is to recover the missing data points leveraging the observed values and intrinsic dependencies within the series. Statistical time series methods, such as autoregressive (AR) models (Meek et al., 2002; Fuller and Hasza, 1981) and autoregressive integrated moving average (ARIMA) models (Newbold, 1983; Nelson, 1998), are capable of imputing missing values; however, they often lack the flexibility required to accurately model the complex characteristics of time series data, such as trends and seasonality. Deep learning methods have also found extensive application in imputation tasks. From the perspective of network architectures, various structures including TCN (Lea et al., 2017), LSTM (Hochreiter and Schmidhuber, 1997), Multi-Layer Perceptron (MLP), and Transformer (Vaswani et al., 2017) have been applied to time series imputation, such as (Du et al., 2023; Zhang et al., 2023; Yıldız et al., 2022; Yuan et al., 2018; Park et al., 2023; Chen et al., 2024). State space models have also been applied to time series imputation tasks, such as (Hu et al., 2024; Dinh et al., 2017). Recently, methods based on generative methods (Tashiro et al., 2021; Alcaraz and Strodthoff, 2023; Yan et al., 2024) have also achieved great success in probabilistic time series imputation.
5.2. Generative Models
Generative models are probabilistic models that learn the original data distribution and generate new samples by sampling from the learned distribution. The goal is to learn the joint probability distribution $p(x)$ of the training data or the posterior distribution $p(x|y)$ and generate new samples that resemble the original data. Traditional generative models include Hidden Markov Models (Huang, 2023; Li et al., 2021a), Bayesian models (Fang et al., 2024; Kocadagli and Asikgil, 2014) and Gaussian process (Corani et al., 2021; Zhang et al., 2021). Researchers have also proposed many deep learning-based generative models, such as Variational Autoencoders (Kingma and Welling, 2014), Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), Normalizing Flow-based Models (Kobyzev et al., 2021), and Denoising Diffusion Probabilistic Models (Ho et al., 2020).
Since the goal of generative models is to model the distribution of unknown points and generate target points by sampling from the learned distribution, they naturally possess the ability to estimate the uncertainty of unknown points. This makes generative models particularly well-suited for probabilistic time series modeling. Therefore, many methods have been proposed which adopts generative models for probabilistic time series modeling, including VAE-based methods (Desai et al., 2021; Li et al., 2021b; Wu et al., 2025a), GAN-based methods (Yoon et al., 2019; Luo et al., 2018), Normalizing-flow based methods (Rasul et al., 2021b; Dai and Chen, 2022) and diffusion-based methods (Rasul et al., 2021a; Shen et al., 2024). State-space models have proved their effectiveness as diffusion backbones (Alcaraz and Strodthoff, 2023; Yan et al., 2024).
6. Conclusion
In this paper, we propose SSD-TS, which explores the potential of linear state-space model (Mamba) as the backbone structure for diffusion models. Based on the requirements of time series modeling task and the characteristics of time series data, we propose a denoising module for diffusion models with Mamba backbone. The proposed module enables bidirectional structural modeling of sequences and effectively leverages both inter-channel and intra-channel dependencies. Experimental results show that, compared to models with structured state-space models, transformers, and convolutional backbones, our method achieves superior performance in time series imputation task. This demonstrates the capability of the linear state-space Mamba model as a backbone for diffusion models. Moreover, our method achieves favorable results in terms of different deterministic and probabilistic metrics, further validating the effectiveness of our design for time series imputation task.
Acknowledgements.
This work was partially supported by National Natural Science Foundation of China (62472174). Jilin Hu is the corresponding author of the work. We gratefully acknowledge the support of the ECNU Multifunctional Platform for Innovation (001) for providing computational resources.